# Galaxy mergers classification using CNNs trained on Sérsic models, residuals and raw images

### v2

@@ Abstract
Galaxy mergers are crucial for understanding galaxy evolution, and with large upcoming datasets, automated methods, such as Convolutional Neural Networks (CNNs), are essential for efficient detection. It is understood that CNNs classify mergers by identifying deviations from the regular, expected shapes of galaxies, particularly faint features, which are indicative of a merger event. In this work, we present a novel investigation of the relative importance of different morphological components: faint residual features and position and spatial structure, in CNN-based merger classification. Using mock images from the IllustrisTNG simulations processed to mimic Hyper Suprime-Cam (HSC) observations, we fit Sérsic profiles to each galaxy and generate three datasets: original images, model images containing only smooth Sérsic profiles, and residual images highlighting faint features after model subtraction. We train three identical CNNs on these datasets: CNN1 on original images, CNN2 on model images, and CNN3 on residual images. CNN1, trained on full images, achieves the highest accuracy of 74%. CNN2, using only shape information including source position, achieves 70%, while CNN3, using only faint residual features, achieves 68%. We find that galaxy merger classification is possible using either faint features or the position and Sérsic profile information present in residual and model images, respectively. Our results demonstrate that not only faint features but also source position information play complementary roles in merger classification. This has important implications for the design and interpretation of machine learning methods for galaxy morphology, particularly in regimes where specific image components may be enhanced or suppressed.
@@
1 Introduction
In the Lambda-cold dark matter cosmology, dark matter (DM) halos form due to gravitational instability, allowing baryonic matter to cool, accumulate, and form galaxies. These galaxies then grow through both accretion within these halos and mergers. Hence galaxy mergers provide crucial insights into how galaxies evolve over time. The merging process proceeds differently depending on the properties of colliding galaxies. Major mergers, where the stellar mass ratio of the merging components is up to 4:1, can combine their structures, resulting in a significant change in morphology if the components collide at the right speed and angle (Bournaud et al., 2005; Lotz et al., 2010). Conspicuous merger features, such as bridges, double nuclei, and tidal features (tails and plumes) of different sizes and shapes may appear (Toomre & Toomre, 1972). This process can feed supermassive black holes, activating galactic nucleus (AGN) activities (Gao et al., 2020; Di Matteo et al., 2005), and trigger star formation (Pearson et al., 2019b; Ellison et al., 2013). When the mass of one of the merging galaxies is significantly larger than the mass of the other one (minor mergers), the merging process is less violent, and the less massive galaxy is usually incorporated into the larger one, leaving the larger galaxy fairly undisturbed.
@@ 1 Introduction
With the advent of upcoming surveys such as Euclid (Laureijs et al., 2011) and Legacy Survey of Space and Time (LSST; Ivezić et al., 2019) that will collect large amount of data it becomes even more important to create algorithms to automate time-consuming, repetitive tasks such as identifying galaxy mergers. The application of machine learning (ML) in the field of astronomy has experienced a significant surge over the past few years. Recent studies have shown that it is possible to use Convolutional Neural Networks (CNN) to address the visual merger classification problem. Ackermann et al. (2018) used transfer learning to fine-tune CNN trained to classify real life objects, using SDDS observations. Pearson et al. (2019a) trained one CNN with observations from SDSS and another with simulated galaxies from EAGLE observations receiving an accuracy of 91.5% and 65.2% respectively. Combination of CNNs and Linear Networks has led to creation of the North Ecliptic Pole merging galaxy catalogue (Pearson et al., 2022) for $z<0.3$. Ćiprijanović et al. (2020) used a CNN to distinguish between mergers and non-mergers in simulated images from Illustris-1 at $z=2$ with 79% accuracy. Bickley et al. (2021) showed that CNNs outperform Gini-M20 and asymmetry methods in post-merger and isolated galaxy classification.
@@
The previous work on creating a CNN classifier for classifying merger galaxies relied on image preprocessing limited to augmentation and normalization (Ackermann et al., 2018; Pearson et al., 2019a, 2022; Ćiprijanović et al., 2020; Bickley et al., 2021). The wide range of characteristic features of mergers and their irregularity, especially subtle peculiarities, may cause these features to be inadequately captured by the network or get mixed up with other galaxy structures. Mantha et al. (2019) explored the use of Sérsic profile subtraction to extract merging features. They developed a tool that extracts and quantifies residual substructures hosted by galaxies. This tool can help to clarify how merger signatures are identified, could enhance efforts to connect observed galaxy morphology to the underlying physical processes of merging, and improve merger detection techniques.
@@
This paper is structured as follows. In Sect. 2 we present the IllustrisTNG simulations data. In Sect. 3 we describe the data preprocessing, the models’ architecture, and define performance metrics used for evaluation. In Sect. 4 we present performance of the models and study the models behaviour. In Sect. 5 we compare the performance of the models in merger classification task. In Sect. 6, we present our conclusions and future work. We assume a cosmology consistent with the Planck Collaboration et al. (2016) results ($\Omega_{\Lambda,0}=0.6911$, $\Omega_{m,0}=0.3089$, $h=0.677$).
2 Data
2.1 Illustris TNG
@@
Images used for this study come from Next Generation Illustris Simulations (IllustrisTNG, or TNG; Nelson et al., 2019), a cosmological gravo-magnetohydrodynamical simulations of formation and evolution of galaxies assuming cosmological parameters consistent with Planck Collaboration et al. (2016). TNG comprises of three runs, TN50 TNG100, and TNG300, with comoving box sizes of 50, 100, and 300 Mpc $h^{-1}$, respectively. All runs follow the same framework. Star formation follows the Springel & Hernquist model Springel & Hernquist (2003). Stars are formed stochastically after gas reaches the density threshold of $n=0.1Hcm^{-3}$. The formation process obeys the Kennicutt-Schmidt relation (Kennicutt Jr, 1998) and Chabrier (2003) initial mass function. Stars evolve and feed back the mass to the interstellar medium via Type Ia and Type II supernovae. Presence of a redshift-dependent, spatially uniform, ionizing UV background radiation field cools metal-enriched gas. The radiation field of nearby AGN activities further adjust the cooling process by superimposing it with the UV background. When a halo is sufficiently massive, a super massive black hole (SMBH) forms, accretes gas from the nearby area and contributes energy back into its surroundings. For a more comprehensive understanding of IllustrisTNG, we direct the reader to the work of Pillepich et al. (2017).
@@
The merger trees were constructed by directly tracking the baryonic content of the subhalos Rodriguez-Gomez et al. (2015). They were later used to create a dataset of merging galaxies by Margalef-Bentabol et al. (2024). They selected galaxies from TNG100 and TNG300. The dataset consists of galaxies in the redshift range $[0.1,1.0)$. Merging galaxies either had a merger event in the past 0.3 Gyr or will have one in the next 0.8 Gyr. Margalef-Bentabol et al. (2024) only consider major mergers where the stellar mass ratios are greater than 4:1. Additionally, they categorize mergers that are 0.8 to 0.1 Gyr before coalescence as pre-mergers, 0.1 Gyr before to 0.1 Gyr after as on-going mergers, and 0.1 to 0.3 Gyr after the coalescence as post-mergers.
2.2 Mock images
@@
Simulated galaxy images are then processed to mimic HSC observations as described in Margalef-Bentabol et al. (2024). The sum of spectral energy distribution of all stellar particles in the image is passed through the filter to create a projected 2D map. Then each image is convolved with HSC i-band PSF. On top of that, Poisson noise that occurs in optical devices is added, and the image is injected into cutouts of real HSC observations. For a more complete description of how the synthetic images were generated, we direct the reader to (Rodriguez-Gomez et al., 2019). A physical size of the images was set to 160 kpc as it is approximately the typical separation between binary mergers based on simulations (Qu et al., 2017; Moreno et al., 2019). Images are 320x320, 192x192, 160x160, and 128x128 pixels for $[0.1,0.31)$, $[0.31,0.52)$, $[0.52,0.76)$, and $[0.76,1.0)$ redshift bins respectively. We randomly divided the images into three sets in the 75:15:10 ratio. The training set, used to fit the parameters, the validation set, providing an evaluation of a model while tuning the model’s hyperparameters, and finally the test set, used for an unbiased evaluation of the final model. Galaxies were divided, so that galaxies belonging to the same merger sequence are included in only one of the three sets. The distribution of galaxies across redshifts can be seen in Fig. 2, and the distribution of galaxies across merging times in Fig. 2.
@@
3 Methodology
3.1 Sérsic profile fitting
We used statmorph (Rodriguez-Gomez et al., 2019) – an open source Python package for calculating morphologies of galaxy images to fit 2D Sérsic profile of each galaxy in the image. First, we create a segmentation map using the SEP Python package - code derived from the Source Extractor (Bertin & Arnouts, 1996) code base. A segmentation map is a map the same size as the original image in which every pixel has an assigned label such that pixels with the same label belong to the same source, and defines the area over which morphological parameters are calculated. Since each pixel value is a combination of background and light from objects, SEP estimates the background level at every point in the image and subtracts it from the original image to accurately measure the flux of the objects. With background-subtracted data SEP performs object detection. We set the detection threshold to a value of $1.5\sigma$ where $\sigma$ is the global background root mean square and minimum number of pixels required for an object is set to 5. Then overlapping but different sources are separated using a deblending algorithm to isolate the galaxy of interest for calculations.
@@ 3.1 Sérsic profile fitting
3.2 Sérsic profile fitting versus masking approaches for feature isolation
@@
3.3 Deep learning
3.3.1 Convolutional Neural Network
A neural network consists of an input layer, multiple hidden layers, and an output layer. The input is processed through a series of calculations in hidden layers. A convolutional neural network (CNN) introduces convolutional layers comprised of learnable filters that convolve across the input image producing feature maps (FM) that capture spatial features. Convolution as an element-wise multiplication is always a linear transformation. To avoid reducing the problem to linear regression in order to learn more complex structures in the data, FMs are passed through an activation function that is nonlinear. Arguably the most common one, the rectified linear unit (ReLU; Nair & Hinton, 2010) function, returns 0 if it receives any negative input, but for any positive input, it returns its value. A pooling layer is used to summarise features that lie in the close regions of a feature map, making the model invariant to changes in positions of features in the input image. It also down-samples the input image, so there are fewer parameters to learn during the training and reduces the computational effort of the process. Batch normalization layers rely on normalising the output of a hidden layer by rescaling and recentring it before feeding it as the input to the next one. A dropout technique causes some number of randomly chosen layer’s outputs to drop out, which means setting the output to 0 during training. It causes the network to never rely on activation of any specific neurons but forces it to learn additional information and helps preventing overfitting.
3.3.2 Training and architecture
@@
The raw images, model images, and residual images are linearly normalized in a range of 0 to 1. These preprocessed images serve as input for all three CNNs. Each image is processed through a series of three convolutional layers (Conv2D) , using 128 kernels of sizes 13×13, 11×11, and 9×9. After each convolutional layer, the ReLU activation function is applied, followed by batch normalization, a 20% dropout, and 2x2 max pooling. After the four convolutional layers, the output is flattened and passed to two fully connected layers (FC) with 512 and 128 neurons, respectively. In these layers, every neuron is connected to all the outputs from the preceding layer. After each fully connected layer, activation, batch normalization, and 20% dropout are applied again. The output of the network is a single neuron activated by a softmax function that provides values representing the likelihood that the image contains a merger. The loss of the network was determined using the binary cross-entropy function and was optimized with the ADAM (Diederik, 2014) algorithm with learning rate $\alpha=3\cdot 10^{-4}$. The CNN has 8 186 113 total trainable parameters. All networks were trained for 50 epochs. All three networks showed signs of overfitting after around 15 epochs, where training accuracy kept increasing while validation accuracy remained relatively still. It means that the problem does not take longer then 50 epochs to train. We used the set of tuned weights and biases for the epoch that provided the highest value of validation accuracy before showing signs of overfitting for classification for the threshold for which precision is equal to recall. A summary of the data pipeline from the creation of the data to the training of the model is shown in Fig. 5, and the summary of CNN architecture in Table 1.
@@
Figure 5: Data pipeline summary. Three datasets were created — raw images, Sérsic profile images, and residual images. Images were cropped to 128 × 128 pixels, linearly normalized, and divided in a 75:15:10 ratio into the training set, the validation set, and the test set.
Table 1: The CNN architecture.
Normalized input images (raw, model, residual), range [0,1]
Conv2D: 128 filters, 13×13 kernel
ReLU, Batch Normalization, Dropout (20%), MaxPooling 2×2
Conv2D: 128 filters, 11×11 kernel
ReLU, Batch Normalization, Dropout (20%), MaxPooling 2×2
Conv2D: 128 filters, 9×9 kernel
ReLU, Batch Normalization, Dropout (20%), MaxPooling 2×2
Flatten
FC1: 512 neurons
ReLU, Batch Normalization, Dropout (20%)
FC2: 128 neurons
ReLU, Batch Normalization, Dropout (20%)
Output layer: 1 neuron
Softmax
Loss: Binary cross-entropy
Optimizer: Adam ($\alpha=3\cdot 10^{-4}$)
Epochs: 50
8,186,113 trainable parameters
3.3.3 Evaluation metrics

### v3

@@ Abstract
Galaxy mergers are crucial for understanding galaxy evolution, and with large upcoming datasets, automated methods, such as Convolutional Neural Networks (CNNs), are essential for efficient detection. It is understood that CNNs classify mergers by identifying deviations from the regular, expected shapes of galaxies, particularly faint features, which are indicative of a merger event. In this work, we present a novel investigation of the relative importance of different morphological components: faint residual features and position and spatial structure, in CNN-based binary classification of galaxies into merger and non‑merger classes. Using mock images from the IllustrisTNG simulations processed to mimic Hyper Suprime-Cam (HSC) observations, we fit Sérsic profiles to each galaxy and generate three datasets: original images, model images containing only smooth Sérsic profiles, and residual images highlighting faint features after model subtraction. We train three identical CNNs on these datasets: CNN1 on original images, CNN2 on model images, and CNN3 on residual images. CNN1, trained on full images, achieves the highest accuracy of 74%. CNN2, using only shape information including source position, achieves 70%, while CNN3, using only faint residual features, achieves 68%. We find that galaxy merger classification is possible using either faint features or the position and Sérsic profile information present in residual and model images, respectively. Our results demonstrate that not only faint features but also source position information play complementary roles in merger classification. This has important implications for the design and interpretation of machine learning methods for galaxy morphology, particularly in regimes where specific image components may be enhanced or suppressed.
@@
I Introduction
In the Lambda-cold dark matter cosmology, dark matter (DM) halos form due to gravitational instability, allowing baryonic matter to cool, accumulate, and form galaxies. These galaxies then grow through both accretion within these halos and mergers. Hence galaxy mergers provide crucial insights into how galaxies evolve over time. The merging process proceeds differently depending on the properties of colliding galaxies. Major mergers, where the stellar mass ratio of the merging components is up to 4:1, can combine their structures, resulting in a significant change in morphology if the components collide at the right speed and angle (Bournaud et al., 2005; Lotz et al., 2010). Conspicuous merger features, such as bridges, double nuclei, and tidal features (tails and plumes) of different sizes and shapes may appear (Toomre and Toomre, 1972). This process can feed supermassive black holes, activating galactic nucleus (AGN) activities (Gao et al., 2020; Di Matteo et al., 2005), and trigger star formation (Pearson et al., 2019a; Ellison et al., 2013). When the mass of one of the merging galaxies is significantly larger than the mass of the other one (minor mergers), the merging process is less violent, and the less massive galaxy is usually incorporated into the larger one, leaving the larger galaxy fairly undisturbed.
@@ I Introduction
With the advent of upcoming surveys such as Euclid (Laureijs et al., 2011) and Legacy Survey of Space and Time (LSST; Ivezić et al., 2019) that will collect large amount of data it becomes even more important to create algorithms to automate time-consuming, repetitive tasks such as identifying galaxy mergers. The application of machine learning (ML) in the field of astronomy has experienced a significant surge over the past few years. Recent studies have shown that it is possible to use Convolutional Neural Networks (CNN) to address the visual merger classification problem. Ackermann et al. (2018) used transfer learning to fine-tune CNN trained to classify real life objects, using SDDS observations. Pearson et al. (2019b) trained one CNN with observations from SDSS and another with simulated galaxies from EAGLE observations receiving an accuracy of 91.5% and 65.2% respectively. Combination of CNNs and Linear Networks has led to creation of the North Ecliptic Pole merging galaxy catalogue (Pearson et al., 2022) for $z<0.3$. Ćiprijanović et al. (2020) used a CNN to distinguish between mergers and non-mergers in simulated images from Illustris-1 at $z=2$ with 79% accuracy. Bickley et al. (2021) showed that CNNs outperform Gini-M20 and asymmetry methods in post-merger and isolated galaxy classification.
@@
The previous work on creating a CNN classifier for classifying merger galaxies relied on image preprocessing limited to augmentation and normalization (Ackermann et al., 2018; Pearson et al., 2019b, 2022; Ćiprijanović et al., 2020; Bickley et al., 2021). The wide range of characteristic features of mergers and their irregularity, especially subtle peculiarities, may cause these features to be inadequately captured by the network or get mixed up with other galaxy structures. Mantha et al. (2019) explored the use of Sérsic profile subtraction to extract merging features. They developed a tool that extracts and quantifies residual substructures hosted by galaxies. This tool can help to clarify how merger signatures are identified, could enhance efforts to connect observed galaxy morphology to the underlying physical processes of merging, and improve merger detection techniques. In particular, a fully automatic merger‑finding method based on the morphology of residual images after Sérsic model subtraction demonstrates that residual‑based diagnostics can effectively isolate disturbed systems in large samples (Hoyos et al., 2012).
@@
This paper is structured as follows. In Sect. II we present the IllustrisTNG simulations data. In Sect. III we describe the data preprocessing, the models’ architecture, and define performance metrics used for evaluation. In Sect. IV we present performance of the models and study the models behaviour. In Sect. V we compare the performance of the models in merger classification task. In Sect. VI, we present our conclusions and future work. We assume a cosmology consistent with the Planck Collaboration et al. (2016) results ($\Omega_{\Lambda,0}=0.6911$, $\Omega_{m,0}=0.3089$, $h=0.677$).
II Data
II.1 Illustris TNG
@@
Images used for this study come from Next Generation Illustris Simulations (IllustrisTNG, or TNG; Nelson et al., 2019), a cosmological gravo-magnetohydrodynamical simulations of formation and evolution of galaxies assuming cosmological parameters consistent with Planck Collaboration et al. (2016). TNG comprises of three runs, TN50 TNG100, and TNG300, with comoving box sizes of 50, 100, and 300 Mpc $h^{-1}$, respectively. All runs follow the same framework. Star formation follows the Springel & Hernquist model Springel and Hernquist (2003). Stars are formed stochastically after gas reaches the density threshold of $n=0.1Hcm^{-3}$. The formation process obeys the Kennicutt-Schmidt relation (Kennicutt Jr, 1998) and Chabrier (2003) initial mass function. Stars evolve and feed back the mass to the interstellar medium via Type Ia and Type II supernovae. Presence of a redshift-dependent, spatially uniform, ionizing UV background radiation field cools metal-enriched gas. The radiation field of nearby AGN activities further adjust the cooling process by superimposing it with the UV background. When a halo is sufficiently massive, a super massive black hole (SMBH) forms, accretes gas from the nearby area and contributes energy back into its surroundings. For a more comprehensive understanding of IllustrisTNG, we direct the reader to the work of Pillepich et al. (2017).
@@
The merger trees were constructed by directly tracking the baryonic content of the subhalos Rodriguez-Gomez et al. (2015). They were later used to create a dataset of merging galaxies by Margalef-Bentabol et al. (2024). They selected galaxies from TNG100 and TNG300. The dataset consists of galaxies in the redshift range $[0.1,1.0)$. Merging galaxies either had a merger event in the past 0.3 Gyr or will have one in the next 0.8 Gyr. Margalef-Bentabol et al. (2024) only consider major mergers where the stellar mass ratios are greater than 4:1. Additionally, they categorize mergers that are -0.8 to -0.1 Gyr before coalescence as pre-mergers, -0.1 Gyr before to 0.1 Gyr after as on-going mergers, and 0.1 to 0.3 Gyr after the coalescence as post-mergers.
II.2 Mock images
@@
Simulated galaxy images are then processed to mimic HSC observations as described in Margalef-Bentabol et al. (2024). The sum of spectral energy distribution of all stellar particles in the image is passed through the filter to create a projected 2D map. Then each image is convolved with HSC i-band PSF. On top of that, Poisson noise that occurs in optical devices is added, and the image is injected into cutouts of real HSC observations. For a more complete description of how the synthetic images were generated, we direct the reader to (Rodriguez-Gomez et al., 2019). A physical size of the images was set to 160 kpc as it is approximately the typical separation between binary mergers based on simulations (Qu et al., 2017; Moreno et al., 2019). Images are 320x320, 192x192, 160x160, and 128x128 pixels for $[0.1,0.31)$, $[0.31,0.52)$, $[0.52,0.76)$, and $[0.76,1.0)$ redshift bins respectively. We randomly divided the images into three sets in the 75:15:10 ratio. The training set, used to fit the parameters, the validation set, providing an evaluation of a model while tuning the model’s hyperparameters, and finally the test set, used for an unbiased evaluation of the final model. Galaxies were divided, so that galaxies belonging to the same merger sequence are included in only one of the three sets. The distribution of galaxies across redshifts can be seen in Fig. 2, and the distribution of galaxies across merging times in Fig. 2. The final mock sample contains 502 152502\,152 galaxies, of which 249 679249\,679 are labelled as mergers and 252 473252\,473 as non-mergers. Among the mergers, 174 689174\,689 are pre-mergers, 34 57434\,574 are ongoing mergers, and 40 41640\,416 are post-mergers. The distribution of systems over these classes is summarized in Table 1.
Mergers and non-mergers 502 152
Mergers 249 679
Non-mergers 252 473
Pre-mergers 174 689
On-going mergers 34 574
Post-mergers 40 416
Table 1: Summary of the galaxy counts in the full sample, including mergers, non-mergers, and the subdivision of mergers into pre-mergers, ongoing mergers, and post-mergers.
@@
III Methodology
III.1 Sérsic profile fitting
We used statmorph (Rodriguez-Gomez et al., 2019) - an open source Python package for calculating morphologies of galaxy images to fit 2D Sérsic profile of each galaxy in the image. First, we create a segmentation map using the SEP Python package (Barbary, 2016) - code derived from the Source Extractor (Bertin and Arnouts, 1996) code base. A segmentation map is a map the same size as the original image in which every pixel has an assigned label such that pixels with the same label belong to the same source, and defines the area over which morphological parameters are calculated. Since each pixel value is a combination of background and light from objects, SEP estimates the background level at every point in the image and subtracts it from the original image to accurately measure the flux of the objects. With background-subtracted data SEP performs object detection. We set the detection threshold to a value of $1.5\sigma$ where $\sigma$ is the global background root mean square and minimum number of pixels required for an object is set to 5. Then overlapping but different sources are separated using a deblending algorithm to isolate the galaxy of interest for calculations.
@@ III.1 Sérsic profile fitting
III.2 Sérsic profile fitting versus masking approaches for feature isolation
@@
III.3 Deep learning
III.3.1 Convolutional neural networks
In this work we use a standard CNN for image‑based binary classification of mergers and non‑mergers. CNNs combine convolutional, pooling, and fully connected layers with non‑linear activations and regularization techniques such as batch normalization and dropout (Nair and Hinton, 2010; LeCun et al., 2002; Ioffe and Szegedy, 2015; Wu and Gu, 2015); a full, detailed description of machine learning is beyond the scope of this paper.
III.3.2 Training and architecture
@@
The raw images, model images, and residual images are linearly normalized in a range of 0 to 1. These preprocessed images serve as input for all three CNNs. The architecture is adapted from the ’Method‑5 (CNN2)’ configuration of Margalef-Bentabol et al. (2024), with a small change in kernel size. The output of the network is a single neuron activated by a sigmoid function that provides values representing the likelihood that the image contains a merger. 20% dropout was used before max-pooling layers (Wu and Gu, 2015). The loss of the network was determined using the binary cross-entropy function and was optimized with the ADAM (Diederik, 2014) algorithm with learning rate $\alpha=3\cdot 10^{-4}$. The CNN has 8 966 273 total trainable parameters. All networks were trained for 50 epochs. All three networks showed signs of overfitting after around 15 epochs, where training accuracy kept increasing while validation accuracy remained relatively still. It means that the problem does not take longer then 50 epochs to train. We used the set of tuned weights and biases for the epoch that provided the highest value of validation accuracy before showing signs of overfitting for classification for the threshold for which precision is equal to recall. A summary of the data pipeline from the creation of the data to the training of the model is shown in Fig. 5, and the summary of CNN architecture in Table 4. The CNNs were implemented using the PyTorch (Paszke et al., 2017) deep‑learning library in Python. Training was performed on RTX 4090 GPU. Training a single CNN for 50 epochs on the full training set required approximately 2 hours, and the total training time for all three models was approximately 6 hours.
@@
Figure 5: Data pipeline summary. Three datasets were created - raw images, Sérsic profile images, and residual images. Images were cropped to 128 × 128 pixels, linearly normalized, and divided in a 75:15:10 ratio into the training set, the validation set, and the test set.
III.3.3 Evaluation metrics